# A Comprehensive Survey on Kolmogorov Arnold Networks (KAN)

### v4

@@ 7.1 Review of Recent Research
Kolmogorov-Arnold Networks (KAN) have achieved significant progress in various fields. Below are some recent important studies:
Fundamental Research and Extensions
Liu et al. (2024) proposed the foundational KAN framework, using learnable univariate activation functions to replace fixed activation functions in traditional neural networks. The study shows that KAN outperforms traditional Multilayer Perceptrons (MLP) in terms of accuracy and interpretability, particularly excelling in data fitting and Partial Differential Equation (PDE) solving tasks [5, 25].
Symbolic Regression
Liu et al. (2024) explored the application of KAN in symbolic regression tasks, demonstrating its advantages in discovering the mathematical relationships and physical laws underlying data. Experimental results indicate that KAN can accurately identify the symbolic formulas generating the data even in high-noise environments [5, 6].
Time Series Prediction
Xu et al. (2024) investigated the application of KAN in time series prediction, proposing two variants: T-KAN and MT-KAN. T-KAN is designed to detect concept drift in time series and explain predictions through symbolic regression of the nonlinear relationships between successive time steps. MT-KAN improves prediction performance by effectively leveraging complex relationships in multivariate time series. Experimental results show that these methods significantly outperform traditional approaches in terms of prediction accuracy and model interpretability [6, 18].
Graph-Structured Data Processing
Liu et al. (2024) introduced GKAN (Graph Kolmogorov-Arnold Networks), extending the principles of KAN to graph-structured data. By introducing learnable univariate functions on graph edges, GKAN innovatively alters the way information is processed in graph structures, replacing traditional convolution architectures. Experimental results indicate that GKAN significantly outperforms traditional GCN models in graph classification and node classification tasks [5, 26].
Wavelet-KAN (Wav-KAN)
Bozorgasl et al. (2024) proposed Wav-KAN, which enhances the interpretability and performance of KAN by incorporating wavelet functions into the KAN framework. Wav-KAN leverages wavelet transforms to capture high-frequency and low-frequency components of the input data, improving classification accuracy, training speed, and robustness, especially in high-dimensional data and complex patterns [26, 21].
KAN 2.0: Kolmogorov-Arnold Networks Meet Science
Recent research by Liu et al. (2024) has led to the development of KAN 2.0, a significant evolution of the Kolmogorov-Arnold Networks. This advancement focuses on integrating KAN with scientific methodologies, offering new functionalities such as MultKAN and the KAN compiler. These tools facilitate the extraction of scientific knowledge from data, making KAN more versatile in identifying relevant features, revealing modular structures, and discovering symbolic formulas[19].
the Robustness of Kolmogorov-Arnold Networks
In another critical development, Alter et al. (2024) explored the robustness of KAN in adversarial scenarios. Their work, titled "On the Robustness of Kolmogorov-Arnold Networks: An Adversarial Perspective," provides a thorough examination of KAN’s resilience against adversarial attacks, particularly in image classification tasks. This study highlights the potential of KAN as a secure and reliable model in environments where robustness is crucial, laying the groundwork for future research in enhancing KAN’s defensive capabilities[27].
@@ 7.2 Improvements and Optimizations
Recent years have seen continuous improvements and optimizations in KAN research. Below are some important advancements:
@@
Afzalaghaei et al. (2024) proposed rKAN, which models using rational function bases. This method employs Padé approximations and rational Jacobi functions, significantly improving KAN’s performance in regression and classification tasks. rKAN enhances model accuracy and computational efficiency through improved activation functions and more efficient parameter update mechanisms [28, 26].
Convolutional KAN (Conv-KAN)
Conv-KAN applies the principles of KAN to Convolutional Neural Networks by introducing learnable univariate functions on convolutional kernels. This approach improves the structure of traditional convolutional layers and excels in image recognition and processing tasks, particularly in handling complex spatial patterns and high-dimensional image data [26, 6].
KAN-Based Adversarial Training (KAT)
Huang et al. (2024) proposed KAT, utilizing the mathematical properties of KAN to enhance the robustness of deep neural networks through adversarial training. Experimental results indicate that KAT significantly improves model robustness against adversarial samples [29, 18].
KAN in Quantum Architecture Search (KANQAS)
Recent studies have explored the application of KAN in Quantum Architecture Search, demonstrating its potential in quantum computing and optimization problems. KANQAS optimizes quantum circuit design, significantly enhancing the efficiency and performance of quantum algorithms [5, 18].
These recent research advancements and optimization methods showcase the broad potential and exceptional performance of KAN in various application fields. Through continuous optimization and expansion, KAN is poised to play an increasingly important role in deep learning and data analysis.
@@ 9 Conclusion
These research directions will not only further KAN’s development but also bring new breakthroughs to the fields of deep learning and data analysis. In summary, as an emerging neural network architecture, KAN exhibits tremendous research and application potential through its theoretical foundation, flexible architecture, and broad application prospects. With continuous optimization and expansion, KAN is expected to play an increasingly important role in future scientific research and practical applications.

### v5

@@ 11 Current Research Progress
Kolmogorov-Arnold Networks (KAN) have achieved significant progress in various fields. Below are some recent important studies. In general, current studies mainly focus on the activation functions.
11.1 Other Variables of KANs
Chebyshev Kolmogorov-Arnold Network
SS et al [48]improve KAN with the Chebyshev polynomials, named as Chebyshev KAN (Polynomial-based Kolmogorov-Arnold Network). Chebyshev polynomials, which is famous of approximation properties, particularly rapid convergence and numerical stability, improve accuracy and efficiency of the KAN. In their experiments, they claim that Chebyshev KAN exceeds MLP in MNIST datasets and shows tremendous potential in nonlinear problems. Based on Chebyshev KAN, Mostajeran and Faroughi [49] propose EPI-cKANs (Elasto-Plasticity Informed Chebyshev-based Kolmogorov-Arnold Network). In their work, they improve traditional elasto-plastic constitutive models by Chebyshev KAN. In their experiments, they test the ability of EPI-cKANs in several mechanical behavior of geotechnical materials tasks. The results show that EPI-cKANs can be an alternative model in non-linear materials behavior various engineering and scientific domains.
Wav-KAN (Wavelet Kolmogorov-Arnold Networks)
Bozorgasl and Chen [50] incorporating wavelet functions into the Kolmogorov-Arnold network structure, enabling the model to capture both high-frequency and low-frequency components of the input data efficiently. Compared with Sql-KAN and traditional MLP, Wav-KAN shows great accuracy, faster training speeds, and increased robustness. They highlight Wav-KAN’s promise in developing interpretable and high-performance networks.
PointNet-KAN
Kashefi [51] introduces PointNet-KAN which replaces the MLP layers in PointNet with KAN. In the experiments, PointNet-KAN shows competitive performance to PointNet in both classification and segmentation tasks. Kashefi claim that the KANs is capable of being applied to advanced point cloud processing architectures.
11.2 Other Application and Study of KANs
Geographic Information Science
Liu et al [52] apply to the problem of baseflow identification. They highlight that KANs enhance the interpretability of machine-learned hydrological models and exceed the state-of-the-art existing semi-empirical models in hydrological processes. Cheon [53] integrates Kolmogorov-Arnold Network (KAN) and Convolutional Neural Network (CNN) models including VGG16 and so forth. After experiments on EuroSAT dataset, Cheon finds that ConvNeXt paired with KAN, named KCN, shows the best performance and might be a promising alternative for efficient image analysis in the RS field.
Computer Vision
Yu et al [54] explore the KANs for image quality assessment by comparing KANs with different activation functions. After rigorous and detailed comparative experiments, they introduce TaylorKAN with mid-level feature input, which shows great performance on 3 realistic image databases (BID2011, CID2013 and CLIVE).
CEST MRI (Chemical Exchange Saturation Transfer Magnetic Resonance Imaging)
Wang et al [55] demonstrate that KAN can be a promising way to analysis CEST MRI for the first time. In their experiment, they find that KAN exceed MLP in extrapolating the CEST fitting parameters although it need more time to train. Moreover, they find that Voxel-wise correlation analysis show that the 4 CEST fitting parameters generated by KAN have a higher Pearson coefficients than the MLP results, which shows the advance of KAN.
11.3 Model Integration and Application Expansion
KAT (Kolmogorov–Arnold Transformer)
Transformer, known as the cornerstone of mordern deep learning, utilizes MLP as the basic block. Yang and Wang [56] replace MLP layers in transformer with KAN, called KAT (Kolmogorov–Arnold Transformer). They claim that their innovation exceed traditional MLP transformer in some vision tasks, including image recognition, object detection, and semantic segmentation.
KAMoE
Based on Gated Residual Kolmogorov-Arnold Networks (GRKAN) Inzirillo and Genet [57] introduce KAMoE, which is a novel framework for Mixture of Experts (MoE). They claim that KAN can be a tool to enhance Multi-task Learning tasks although its standalone is still under discussion.
11.4 Robustness and Generalization
Alter, Lapid and Sipper[alter2024robustnesskolmogorovarnoldnetworksadversarial] test the robustness of KAN under adversarial conditions, especially, they put emphasis on image classification tasks. After experiments, they claim that KANs may be particularly effective in scenarios where larger models are feasible and robustness is critical. However, after comparing the generalization of MLP and KAN, they highlight that KAN has low transferability of information, which show the different structure between KAN and MLP.
11.5 Theoretical Analysis
Gao and Tan [58] studied gradient descent (GD) and stochastic gradient descent (SGD) of KAN. They claim that using SGD to optimize KAN can be used to enhance the performance on some machine learning tasks. Moreover, they analysis the global convergence of GD and SGD for physics-informed KANs, which might be used to optimize physics-informed KANs. Wang et al [59] carefully compared KAN and MLP from a theoretical perspective. From the perspective of approximation, MLP and KAN is very similar sharing parameterize similar classes of functions. However, they find that KANs lack the spectral bias toward low frequencies that is characteristic of MLPs. They claim that their research results can contribute to computing applications like solving PDEs.
@@ 11.6 Improvements and Optimizations
Afzalaghaei et al [60] proposed rKAN, which models using rational function bases. This method employs Padé approximations and rational Jacobi functions, significantly improving KAN’s performance in regression and classification tasks. rKAN enhances model accuracy and computational efficiency through improved activation functions and more efficient parameter update mechanisms.
ReLU-KAN
Qiu et al [61] replace basis spline functions (B-Spline function) with singly matrix addition, dot production and ReLU activation functions. They claim that their new structure of KAN can help handle the problem of catastrophic forgetting and its backpropagation is much faster efficient GPU utilization. Also, they mention that ReLU-KAN demonstrates enhanced fitting performance for complex problems, which exceed traditional KAN.
Kolmogorov-Arnold Networks Using Sinusoidal Activation Functions (SineKAN)
Reinhardt, Dinesh and Gleyzer [62] utilize sinusoidal activation functions as B-Spline function. In their experiments, they compare SineKAN with B-SplineKAN on MNIST dataset. They claim that The SineKAN model outperforms B-SplineKAN in performance, scalability, and speed, achieving up to 9x faster benchmarks.
Unbound Kolmogorov-Arnold Network (UKAN)
Moradzadeh et al [63] introduce the Unbound Kolmogorov-Arnold Network (UKAN), which combines MLP layers with KANs along with an efficient GPU implementation. They claim that UKAN can reduce time complexity. In their experiments, they test their model on Moon and MNIST datasets. The results show that UKAN get almost 100 percent accuracy.
Higher-order-ReLU-KANs (HRKANs)
So and Yung [64] introduce HRKANs, which uses a higher-order-ReLU as the activation function. In their experiments, they exam URKANs on two famous and representative PDEs linear Poisson equation and nonlinear Burgers’ equation with viscosity. They claim that URKANs achieve superior fitting accuracy, enhanced robustness, and faster convergence.
Selectable KAN (S-KAN)
Yang et al [65] introduce S-KAN, which can select activation functions in the pool. Also, they expand S-KAN and propose an activation space selectable Convolutional KAN (S-ConvKAN). They claim that the performance of S-KAN is significantly enhanced by using multiple different learnable nonlinear activations. Also, they find that S-COnvKAN is suitable for general computer vision image tasks.
@@ 13 Conclusion
These research directions will not only further KAN’s development but also bring new breakthroughs to the fields of deep learning and data analysis. In summary, as an emerging neural network architecture, KAN exhibits tremendous research and application potential through its theoretical foundation, flexible architecture, and broad application prospects. With continuous optimization and expansion, KAN is expected to play an increasingly important role in future scientific research and practical applications. .